# Uncertainty-enabled machine learning for emulation of regional sea-level change caused by the Antarctic Ice Sheet

### v1

@@ 2.3 Partitioning of training, validation, and test datasets
The entire dataset of ISMIP6 AIS mass change fields, along with their corresponding numerically solved sea-level change outputs, was split into training, validation, and test datasets, as is commonly done in ML (Hastie et al., (2009)), with 80% used for training, 10% used for validation, and 10% used for a hold-out test set. As mentioned above, this constitutes 1,207 pairs of AIS mass change and global relative sea-level change fields across the different ice-sheet models, experiments, and time levels.
In order to ensure representation of each year across the three data splits, the input-output pairs from the 71 models per each year were randomly selected to be in the training, validation, and test sets. That is, for a given year, 57 were randomly selected to be in the training set, 7 were randomly selected to be in the validation set, and 7 were randomly selected to be in the test set. The final split of data consisted of 969 in training, 119 in validation, and 119 in testing, where the inputs of each split were standardized to have 0 mean and unit variance. Before applying a variety of ML models (described in Section 3) to this dataset, we preprocessed the sea-level dataset by removing those elements that were zeros throughout the entire dataset (i.e., cells without grounded ice at any time, almost two-thirds of the grid cells, Fig. 1a,b). After preprocessing, the length of the flattened input vector was 205,928, and the output vector was the 27 long-term tide gauge locations.
@@ 3 ML models
Our overarching objective is to construct a machine learning (ML) model that takes as input a vector of ice thickness changes across Antarctica derived from ISMIP-6 simulations and outputs a prediction of the simulation of resultant RSL change at 27 long-term tide gauge locations (Fig. 1c,d). That is, in our RSL change emulator, the input is the flattened vector of ice-thickness changes, and the output is a prediction of the resultant RSL change for the 27 long-term tide gauge locations. Ideally, the ML model should accurately predict simulated RSL change values, be computationally more affordable than running the full numerical simulation, and also enable statistically valid uncertainty quantification of predictions. In what follows, we describe a variety of ML implementations that are constructed and compared: a feedforward neural network (NN), a conditional variational autoencoder (CVAE), a Gaussian process (GP) acting on a reduced-dimensional space, and a random forest (RF) acting on a reduced-dimensional space. Our main finding is that the NN, CVAE, and GP models yield similar levels of accuracy for this application, but the NN and CVAE yield the best uncertainty quantification in terms of prediction intervals using a simple linear regression postprocessing technique; moreover, prediction intervals are of a useful length. In what follows, we describe these ML approaches and our implementations in more detail.
@@ 3.1 Feedforward neural network (NN)
A feedforward neural network (NN) is a fundamental component of deep neural networks. The primary objective of an NN is to learn (or approximate) a mapping $\bm{y}=\bm{f}(\bm{x};\bm{\theta})$ from input $\bm{x}$ to output $\bm{y}$, accomplished through the composition of functions governed by parameters $\bm{\theta}$ that consist of weights and bias terms. The approximation task is equivalent to determining the best estimate $\hat{\bm{\theta}}$ that minimizes a loss function (e.g., mean-squared error) using a training set, which is typically updated iteratively with backpropagation (Rumelhart et al.,, 1986).
It is common to use well-known optimization routines such as stochastic gradient descent (SGD) (Bottou,, 2010) and ADAM (Kingma and Ba,, 2017) for fitting neural networks. A feedforward neural network (NN) consists of an input layer, $L$ hidden layers, and an output layer, where $L$ denotes the number of hidden layers. $\ell^{th}$ hidden layer is equipped with $d_{\ell}$ neurons, which serve as the dimensions of hidden states.
The layers’ units (neurons) are interconnected through a linear transformation and an activation function, denoted as $g(\bm{x}^{\top}\bm{w}+\bm{b}$), where $\bm{w}$, $\bm{b}$, and $g(\cdot)$ correspond to the weight, bias, and an activation function, respectively. A sequence of these operations across each layer can yield a high-quality approximation to complex functions. The potential for such approximation is enhanced by the activation function, such as the rectified linear unit (RELU), which introduces non-linearity. In addition to its rich approximating capabilities, user-friendly software, such as PyTorch by Paszke et al., (2019), has contributed to the popularity of NNs. For a more detailed introduction to NNs, refer to Goodfellow et al., (2016).
In our implementation, we utilize a neural network architecture comprising $L=4$ hidden layers. We use the RELU activation function between layers, while a linear activation function is applied between the last hidden layer and the output layer. The number of neurons within each layer is $(d_{1},d_{2},d_{3},d_{4})=(3200,1600,800,400)$. Using root mean squared error (RMSE) as the loss function, we train the model with the SGD optimizer with a learning rate of 0.001. Throughout the training process, we monitor the convergence of the loss on the validation set. To prevent overfitting, we also introduce dropout (Srivastava et al.,, 2014) with a probability 0.3 between layers. We implement the NN using Pytorch version 2.0.1 with Python version 3.9.12.
@@ 3.2 Conditional variational autoencoder (CVAE)
The CVAE is a widely-used modern neural network architecture that couples an encoder to a decoder, with the intermediate layer between them representing a latent, lower-dimensional space. The CVAE is appealing and can be effective in constructing generative models (e.g., a molecular generative model (Lim et al.,, 2018) and zero-shot learning generative model (Mishra et al.,, 2018)). We consider a CVAE in addition to the feed-forward NN in order to attempt a generative approach for the sea-level change emulator, since by sampling from the appropriate latent space, one can generate samples of predictions for a given input as opposed to a single prediction; indeed this approach serves as a precursor for generative artificial intelligence (AI) approaches such as diffusion models (Ho et al.,, 2020).
@@
The origin of the CVAE can be traced back to the autoencoder (AE) (Kramer,, 1991) and the variational autoencoder (VAE) (Kingma and Welling,, 2013). In general, an AE is trained to reproduce the original input, thereby learning a reduced-dimension latent representation that captures much of the information of the higher dimensional input data. Although the architecture of the VAE is similar to the AE, the VAE assumes that the reduced-dimensional latent space follows a certain distribution, often a multivariate normal with 0 mean and identity covariance matrix. The key advantage of the VAE over the AE is that it can serve as a generative model by drawing a sample from the distribution of the latent, intermediate layer.
@@
While both the CVAE and the VAE are similar, the CVAE possesses an additional feature: the ability to instruct the model to predict the output based on conditional information. In essence, a prediction for a new input is determined by concatenating a random draw from the latent space with conditional information, which in our setting is the vector of ice-thickness changes. This concatenated vector, which combines conditional information with latent information, is passed through the decoder to make a prediction. For a visual representation, see Figure 2; more details are described in the tutorial of Doersch, (2016).
Figure 2: CVAE example with an intermediate layer that follows a multivariate normal distribution. Note that conditional information is provided in the decoder.
@@
Our implementation of the CVAE for the sea-level change emulator takes the entire vector of ice-thickness changes as conditional information, concatenates it with a latent draw, and passes the concatenated vector through the decoder to make a prediction. Specifically, the encoder comprises $L=4$ hidden layers with neurons ($d_{1}$, $d_{2}$, $d_{3}$, $d_{4}$) = (3200, 1600, 800, 400), and it outputs 200 pairs of parameters ($\mu_{i}$, $log\sigma^{2}_{i}$), where $i=1,...,200$. These parameters correspond to the parameters of the posterior intermediate layer distribution $P(\bm{Z}|\bm{X})$, which is assumed to follow a multivariate normal distribution with a mean vector $\bm{\mu}=(\mu_{1},...,\mu_{200})$ and a diagonal covariance matrix $diag(\sigma^{2}_{1},...,\sigma^{2}_{200})$. Then, the decoder takes as input the concatenated vector of the ice-sheet thickness vector (i.e., the conditional information) and the transformed samples from the prior distribution $P(\bm{Z})$ for the intermediate layer. Assuming $P(\bm{Z})$ follows a multivariate normal distribution with 0 mean and identity covariance matrix, the transformation is achieved through re-parametrization $Z_{i}|\bm{X}=\mu_{i}+\sigma_{i}Z_{i}$. Similar to the encoder, the decoder also consists of $L=4$ hidden layers with the same number of neurons as the encoder. It outputs the predicted sea-level change at 27 long-term tide gauge locations. Because the CVAE is predicting sea-level change values as opposed to reconstructing its own input, this is a modification of the usual AE approach.
@@
3.3 Baseline comparisons using Gaussian process and random forest
We use baseline comparison emulators to assess the performance of the NN and CVAE emulators. These baseline comparisons are a stationary Gaussian process with a squared-exponential covariance function (GP) and a random forest (RF). Gaussian processes have been a common choice to build emulators for physical applications (Higdon et al.,, 2008; Gu et al.,, 2018) and hence serve as a natural comparison to NN and CVAE emulators. They are primarily used because of closed-form multivariate normal distributions conditioning on training data; see Gramacy, (2020) for exact formulations.
For the baseline GP emulator, we use a GP that maps from a reduced dimensional input space to the sea-level change at each of the 27 long-term tide gauge locations. The dimension-reduction technique we adopt in this work is discussed subsequently. The GP uses a squared-exponential kernel (Rasmussen,, 2003; Gramacy,, 2020) whose parameters, including a nugget term, are fit with a statistical technique known as profile likelihood maximization, the mathematical steps of which are described in Chapter 5 of Gramacy, (2020). Our implementation of profile likelihood was coded in the R programming language. While there exist some recent variants of GPs such as deep Gaussian processes (e.g., Sauer et al., (2023)) that are capable of handling nonstationarity, we use a stationary GP with a squared-exponential covariance function only to serve as a baseline comparison.
RFs have been widely adopted as off-the-shelf ML models, making an RF model an ideal additional baseline comparison (e.g., see Hooten et al., (2011) and Gopalan and Wikle, (2022) for emulators of physical processes based on RFs). A detailed introduction to RF models is provided in Hastie et al., (2009) and James et al., (2013). For baseline comparisons, we use the “RandomForestRegressor" function in a package “scikit-learn" (Pedregosa et al.,, 2011) to train an RF in Python 3.9.17. We use 100 decision trees (default) and mean squared error (MSE) as the loss function for this baseline RF. As is the case for the GP baseline, an RF is trained for each of the 27 long-term tide gauge locations using the reduced dimensional input.
While the NN and CVAE models take as input an entire vector of ice thickness changes across Antarctica and perform dimension reduction within their respective architectures, the GP and RF models use a reduced-dimension space as input. Dimension reduction was performed by projecting input vectors onto a reduced-rank principal-component basis; 250 principal components were used, which represent almost 98.7% of the total variation of the training set inputs. This technique is standard for emulating high-dimensional computer simulator output (e.g., Higdon et al., (2008), Hooten et al., (2011), Gopalan and Wikle, (2022)), but in this work we use dimension reduction on the space of inputs for the GP and RF emulators, which is likely required as the dimension of inputs grows substantially – the input vector length is on the order of 100,000 whereas there are on the order of 1000 data vectors. For instance, computing the distance matrix needed for a GP model is an $O(n^{2}m)$ operation when there are $n$ length $m$ vectors, which for $m>>n$ can be more substantial than the $O(n^{3})$ inverse and determinant calculations needed for GPs.
@@ 4 Uncertainty quantification
We construct prediction intervals as a means of quantifying uncertainty in emulated sea-level change values at the 27 long-term tide gauge locations. For the NN, CVAE, and RF emulators, we use a simple linear regression postprocessing technique, discussed in what follows, in order to generate prediction intervals. For GPs, the distribution of predictions given training points is multivariate normal with exact mean and covariance (Rasmussen, (2003); Gramacy, (2020)). Prediction intervals can thus be derived for the GP emulator by using the appropriate quantiles of the resultant Gaussian distribution. Before explaining the linear regression technique, we note that we found the empirical prediction intervals solely constructed with Monte Carlo dropout for both the NN and CVAE were far too narrow; a similar phenomenon was exhibited with generated samples from the CVAE.
As reviewed in Lovegrove and Siegert, (2023), linear regression methods have been used to postprocess outputs from numerical climate models by using the numerical model output as an input variable (i.e., covariate) for either a simple or multiple linear regression, the latter case when there are multiple covariates. In the weather forecasting literature, this has been referred to as model output statistics (Glahn and Lowry,, 1972; Glahn et al.,, 2009). Lovegrove and Siegert, (2023) show with a global temperature forecasting example that fitting simple linear regressions yields prediction intervals for temperature that have close to nominal coverage probabilities; i.e., a (1-$\alpha$)prediction interval as derived from the simple linear regression covers the true value with probability near (1-$\alpha$)%. In addition, such an approach has been used for calibrating probabilities generated from ML models for categories via logistic regression, referred to as Platt scaling (Platt et al.,, 1999).
Our approach follows this idea: for a given location, we regress the simulated sea-level change values onto the ML predictions for all of the simulation data in the validation set using a simple linear regression. This is also similar to the split-conformal method of Lei et al., (2018) because we use a dataset split from the original training data to fit the linear regression models, as to avoid overfitting by training both the ML model and the simple linear model on the same data. However, we are able to obtain conditional coverage by fitting different linear regressions for each location. Moreover, additional covariates could be included into the linear regression to improve conditional calibration. Additionally, a per-location bias is accounted for by the intercept term of the simple linear regression models.

### v2

@@ 2.3 Partitioning of training, validation, and test datasets
The entire dataset of ISMIP6 AIS mass change fields, along with their corresponding numerically solved sea-level change outputs, was split into training, validation, and test datasets, as is commonly done in ML (hastie2009elements), with 80% used for training, 10% used for validation, and 10% used for a hold-out test set. As mentioned above, this constitutes 1,207 pairs of AIS mass change and global relative sea-level change fields across the different ice-sheet models, experiments, and time levels.
In order to ensure representation of each year across the three data splits, the input-output data pairs from the 71 ice-sheet simulations per each year were randomly selected to be in the training, validation, and test sets; that is, we stratify by year when constructing the training, validation, and test splits. Of the 71 total ice-sheet simulations for a given year, 57 were randomly selected to be in the training set, 7 were randomly selected to be in the validation set, and 7 were randomly selected to be in the test set, yielding 80% for training, 10% for validation, and 10 % for test. The final split of data consisted of 969 in training, 119 in validation, and 119 in testing, where the inputs of each split were standardized to have 0 mean and unit variance. The length of the flattened input vector for each time step of each simulation was 579,121, and the output vector was the 27 long-term tide gauge locations.
@@ 3 ML models
We constructed machine learning (ML) models that take as input a vector of ice thickness changes across Antarctica derived from ISMIP-6 simulations and outputs a prediction of the simulation of resultant RSL change at 27 long-term tide gauge locations (Fig. 1c,d). That is, in our RSL change emulator, the input is the flattened vector of ice-thickness changes, and the output is a prediction of the resultant RSL change for the 27 long-term tide gauge locations. Ideally, the ML model should accurately predict simulated RSL change values, be computationally more affordable than running the full numerical simulation, and also enable statistically valid uncertainty quantification of predictions. In what follows, we describe neural network ML implementations that are constructed and compared: a feedforward neural network (NN) and a conditional variational autoencoder (CVAE). We choose neural networks due to their prevalence in many widely-used modern ML implementations, including transformers in large language and vision models (vaswani2017attention), for physics-informed ML (raissi2019physics), and in generative artificial intelligence (achiam2023gpt). We discuss other possibilities at the end of the section.
@@ 3.1 Feedforward neural network (NN)
A feedforward neural network (NN) is a fundamental component of deep neural networks. The primary objective of an NN is to learn (or approximate) a mapping $\bm{y}=\bm{f}(\bm{x};\bm{\theta})$ from input $\bm{x}$ to output $\bm{y}$, accomplished through the composition of functions governed by parameters $\bm{\theta}$ that consist of weights and bias terms. The approximation task is equivalent to determining the best estimate $\hat{\bm{\theta}}$ that minimizes a loss function (e.g., mean-squared error) using a training set, which is typically updated iteratively with backpropagation (Rumelhart).
It is common to use well-known optimization routines such as stochastic gradient descent (SGD) (sgd) and ADAM (adam) for fitting neural networks. A feedforward neural network (NN) consists of an input layer, $L$ hidden layers, and an output layer, where $L$ denotes the number of hidden layers. $\ell^{th}$ hidden layer is equipped with $d_{\ell}$ neurons, which serve as the dimensions of hidden states.
The layers’ units (neurons) are interconnected through a linear transformation and an activation function, denoted as $g(\bm{x}^{\top}\bm{w}+\bm{b}$), where $\bm{w}$, $\bm{b}$, and $g(\cdot)$ correspond to the weight, bias, and an activation function, respectively. A sequence of these operations across each layer can yield a high-quality approximation of complex functions. The potential for such approximation is enhanced by the activation function, such as the rectified linear unit (RELU), which introduces non-linearity. In addition to its rich approximating capabilities, user-friendly software, such as PyTorch by pytorch, has contributed to the popularity of NNs. For a more detailed introduction to NNs, refer to goodfellow.
In our implementation, we utilize a neural network architecture comprising $L=4$ hidden layers. Due to the multiple layers, we consider this to be a deep neural network (goodfellow). We use the RELU activation function between layers, while a linear activation function is applied between the last hidden layer and the output layer. The number of neurons within each layer is $(d_{1},d_{2},d_{3},d_{4})=(3200,1600,800,400)$. Using root mean squared error (RMSE) as the loss function, we train the model with the SGD optimizer with a learning rate of 0.001. Throughout the training process, we monitored the convergence of the loss on the validation set. To prevent overfitting, we also introduce dropout (dropout) with a probability 0.3 between layers. We implement the NN using Pytorch version 2.0.1 with Python version 3.9.12.
@@ 3.2 Conditional variational autoencoder (CVAE)
Figure 2: Illustration of the CVAE architecture. The encoder, in the left-most box, is a feed-forward neural network that maps to the latent space in the middle of the diagram. This latent space is lower dimensional than the input space. For generating output, a latent multivariate normal vector, depicted with a density plot, is sampled and concatenated with conditional information, and then it is subsequently passed through the decoded network to produce a prediction. Note that this a generative model as multiple outputs can be generated with the same conditional information so long as new latent vectors are sampled.
@@
The CVAE is a widely-used modern neural network architecture that couples an encoder to a decoder, with the intermediate layer between them representing a latent, lower-dimensional space. The CVAE is appealing and can be effective in constructing generative models (e.g., a molecular generative model (lim) and zero-shot learning generative model (Mishra_2018_CVPR_Workshops)). We consider a CVAE in addition to the feed-forward NN in order to attempt a generative approach for the sea-level change emulator, since by sampling from the appropriate latent space, one can generate samples of predictions for a given input as opposed to a single prediction; indeed this approach serves as a precursor for generative artificial intelligence (AI) approaches such as diffusion models (ho2020denoising).
@@
The origin of the CVAE can be traced back to the autoencoder (AE) (autoencoder) and the variational autoencoder (VAE) (vae). In general, an AE is trained to reproduce the original input, thereby learning a reduced-dimension latent representation that captures much of the information of the higher dimensional input data. Although the architecture of the VAE is similar to the AE, the VAE assumes that the reduced-dimensional latent space follows a certain distribution, often a multivariate normal with 0 mean and identity covariance matrix. The key advantage of the VAE over the AE is that it can serve as a generative model by drawing a sample from the distribution of the latent, intermediate layer.
@@
While both the CVAE and the VAE are similar, the CVAE possesses an additional feature: the ability to instruct the model to predict the output based on conditional information. In essence, a prediction for a new input is determined by concatenating a random draw from the latent space with conditional information, which in our setting is the vector of ice-thickness changes. This concatenated vector, which combines conditional information with latent information, is passed through the decoder to make a prediction. For a visual representation, see Figure 2; more details are described in the tutorial of doersch2016tutorial.
@@
Our implementation of the CVAE for the sea-level change emulator takes the entire vector of ice-thickness changes as conditional information, concatenates it with a latent draw, and passes the concatenated vector through the decoder to make a prediction. Specifically, the encoder comprises $L=4$ hidden layers with neurons ($d_{1}$, $d_{2}$, $d_{3}$, $d_{4}$) = (3200, 1600, 800, 400), and it outputs 200 pairs of parameters ($\mu_{i}$, $log\sigma^{2}_{i}$), where i=1,,200i=1,...,200. These parameters correspond to the parameters of the posterior intermediate layer distribution $P(\bm{Z}|\bm{X})$, which is assumed to follow a multivariate normal distribution with a mean vector $\bm{\mu}=(\mu_{1},...,\mu_{200})$ and a diagonal covariance matrix $diag(\sigma^{2}_{1},...,\sigma^{2}_{200})$. Then, the decoder takes as input the concatenated vector of the ice-sheet thickness vector (i.e., the conditional information) and the transformed samples from the prior distribution $P(\bm{Z})$ for the intermediate layer. Assuming $P(\bm{Z})$ follows a multivariate normal distribution with 0 mean and identity covariance matrix, the transformation is achieved through re-parametrization $Z_{i}|\bm{X}=\mu_{i}+\sigma_{i}Z_{i}$ (vae). Similar to the encoder, the decoder also consists of $L=4$ hidden layers with the same number of neurons as the encoder. It outputs the predicted sea-level change at 27 long-term tide gauge locations. Because the CVAE is predicting sea-level change values as opposed to reconstructing its own input, this is a modification of the usual AE approach.
@@
3.3 Other ML options
We considered additional ML models besides neural networks as baseline comparisons. These baseline comparisons are a stationary Gaussian process with a squared-exponential covariance function (GP) and a random forest (RF). In our preliminary work, we found that the neural-network emulators performed just as accurately or more accurately than the baselines aforementioned. Additionally, the Gaussian process and random-forest emulators used an additional dimensionality reduction step to allow for scalability, whereas the neural networks are able to handle the order 100,000 dimensional input without a preliminary dimension reduction procedure. The use of dimensionality reduction can be accomplished in a number of ways (e.g., principal components as we implemented or random subspace projections), and so an additional dimension-reduction step introduces a conflating factor in evaluating performance of the GP and RF models. These considerations, coupled with the fact that neural networks are the predominant ML model of choice for modern AI applications, led us to proceed with only the neural-network emulators in comparing to the physics-based kernel sensitivity method, which is next described. Nonetheless, a description and review of the GP and RF approaches is presented in the Appendix.
4 Kernel sensitivity method
Although applications of the sea-level equation have been long-established for predicting RSL caused by a specified change in land-ice thickness (e.g. Farrell1976; Kendall2005; Mitrovica2003), such applications require re-solving the sea-level equation for each unique spatial pattern of land-ice mass change. In contrast, in recent years, multiple formulations have been made that invert the relationship between land-ice mass change and RSL to produce sensitivity maps of RSL at a given location to the spatial patterns of ice thickness change (Larour2017; Crawford2018; Mitrovica2018; Al-Attar2024). These sensitivity kernel approaches are attractive for the problem addressed in this study of evaluating many realizations of the impact of ice-sheet thickness change on RSL at a modest number of coastal locations; by calculating the sensitivity kernels once for each coastal site of interest, the RSL at each site due to any pattern of ice-sheet thickness changes can subsequently be evaluated for the computational cost of a single vector multiplication.
Of the RSL sensitivity-kernel methods that have been described, we selected the recent method of Al-Attar2024 for this application. It is restricted to the simpler elastic fingerprint problem we consider here, ignoring the complexities of viscoelastic sensitivity kernels (c.f. Crawford2018). It includes the rotational feedbacks that are important for AIS mass changes (Cederberg2023; Roffman2023). Lastly, it is easy to implement with any elastic fingerprint model and does not require automatic differentiation or specialized model implementation.
Although the sensitivity-kernel method is attractive for our application, it includes some simplifications relative to the full sea-level model used. In addition to the limitation of elastic rheology, the Al-Attar2024 method has the restriction of requiring a fixed shoreline position, which renders the elastic fingerprint equation self-adjoint and thus possible to generate sensitivity kernels with a standard forward model. For the sea-level changes ($\mathcal{O}(10^{-2})$ m) from the ISMIP6 AIS ensemble for 2015-2100, shoreline migration from inundation of coastal areas is expected to be negligible. However, shoreline migration due to the retreat of grounded marine ice sheets will not be, and must be treated approximately in a fixed shoreline model. A difference between the Han2022 and Al-Attar2024 models is that the Han2022 model considers the full ice thickness of marine ice sheets and iteratively evaluates the flotation condition as part of the RSL solution. Additionally, it has an evolving ocean mask that adjusts as grounded ice retreats from or advances into a grid cell, allowing grid cells to dynamically switch between being occupied by ice or ocean (as described in Kendall2005).
We apply the sensitivity kernel method of Al-Attar et al. (2024) to the sea-level code of Al-Attar (2023), since it is not compatible with the evolving shoreline capability of the Han model. We verified the code in two steps: 1. Using the same model parameters and topography and ice-sheet thickness changes, we confirmed that the difference in RSL between the models of Han2022 and Al-Attar_2023 are small ($1\times 10^{-3}$ m) for forward model solutions; 2. We confirmed that the difference in RSL between the forward model solution and sensitivity kernel method with the Al-Attar_2023 code are very small ($1\times 10^{-4}$ m), as reported by Al-Attar2024. Although we are not further developing the sensitivity kernel method in this work, this is the first systematic application of it that we are aware of to future projections of sea-level change caused by the Antarctic Ice Sheet.
To generate sensitivity kernels, we followed the procedure of Al-Attar2024. First, we prepared a reference dataset of initial topography and ice thickness. Because each ISMIP6 modeling group was free to adopt different initial conditions, we calculated the average topography and ice thickness across all models. Additionally, we calculated an ice mask for the kernel method that represents the union of the initial grounded ice across all ISMIP6 models; this ensured that changes in grounded-ice thickness from any of the ISMIP6 models were treated as a load change when applied to the kernel method, while producing an ocean area approximately appropriate for the initial ice extent from all models. With these fields, we calculated sensitivity kernels for each of the 27 coastal cities of interest, by applying a point load smoothed over a characteristic length-scale of 1 degree as described by Al-Attar2024. To subsequently apply the sensitivity kernels, we integrated the sensitivity kernel for each city against specified ice thickness changes. To account for the fixed shoreline assumption, we restricted ice thickness change to changes in height above flotation, a difference from how ice thickness change was applied to the Han2022 model. Note that although the theory of the sensitivity kernel method demonstrates it will be of equal accuracy to a fixed-shoreline elastic fingerprint solution (Al-Attar2024), the series of small adjustments from our more complex application of the model of Han2022 (fixed shoreline, restricting load change to height above flotation, assumption of single initial condition and ocean area) introduced error, which in part motivates the comparison against the ML methods described above.
@@ 5 Uncertainty quantification
We construct prediction intervals for simulated sea-level change values as a means of quantifying emulator uncertainty at the 27 long-term tide gauge locations. We considered three main ways of quantifying uncertainty in this study: i) linear regression post-processing, ii) split-conformal inference, and iii) Monte Carlo dropout with latent variable sampling. We describe in more detail these three avenues for uncertainty quantification (UQ) below, and we emphasize that the use of UQ for the sensitivity kernel approach is a novel contribution to the literature.
5.1 Linear regression prediction intervals
As reviewed in lovegrove2023improving, linear regression methods have been used to postprocess outputs from numerical forecasts by using the numerical model output as an input variable (i.e., covariate) for either a simple or multiple linear regression, the latter case when there are multiple covariates. In the weather forecasting literature, this has been referred to as model output statistics (glahn1972use; glahn2009mos). lovegrove2023improving show with a global temperature forecasting example that fitting simple linear regressions yields prediction intervals for temperature that have close to nominal coverage probabilities; i.e., a (1-$\alpha$) prediction interval as derived from the simple linear regression covers the true value with probability near (1-$\alpha$) where $\alpha$ is the miscoverage rate. In addition, such an approach has been used for calibrating probabilities generated from ML models for categories via logistic regression, referred to as Platt scaling (platt1999probabilistic).
We implement this linear regression approach as follows: for a given location, we regress the simulated sea-level change values onto the emulator predictions for all of the simulation data in the validation set using a simple linear regression. This is also similar to the split-conformal method of lei2018distribution because we use a dataset split from the original training data to fit the linear regression models, as to avoid overfitting by training both the emulator model and the simple linear model on the same data. However, we are able to improve upon conditional coverage by fitting different linear regressions for each location. Moreover, additional covariates could be included into the linear regression to improve conditional calibration. Additionally, a per-location bias is accounted for by the intercept term of the simple linear regression models.